A guitar string has an overall length of 1.19 m and a total mass of 2.3 g before being strung on a guitar. Once it is used on the guitar, there is a distance of 70 cm between fixed end points. The guitar string is tightened to a tension of 390 N.

What is the frequency of the fundamental wave on the guitar string?
Answer #1

Length of wire=1.19 m

Mass of wire=2.3 g=0.0023 kg

Mass per unit=0.0023/1.19=0.0019 kg/m

Tension in wire=390 N

Distance between fixed ends=70 cm=0.7 m

Let frequency be f

f=sqrt(390/0.0019)/2*0.7=320.86 Hz

This is the required answer